Scenario 1 for questions 11 and 12A monopolist faces the following demand curve, marginal revenue curve and total cost curve for its product:Q = 400 -2PMR = 200 -QTC = 10Q11.Refer to scenario 1. How much profit does the monopolist earn?


1
Expert's answer
2021-06-08T12:23:06-0400

Solution


Q=400-2P

MR=200-Q

TC=10Q

MR=0

MR=200-Q=0

Q=0

Q is the revenue maximizing output

Q=400-2P

200=400-2P

2P=400-200

P=100 - profit maximizing price

"TR=P\times Q"

=100"\times200=20000"

TC= 10Q

"=10\times200=2000"

Profit= TR-TC

=20000-2000

=18,000 is the profit earned by the monopolist